# Distributionally and Adversarially Robust Logistic Regression via Intersecting Wasserstein Balls

### v1

@@ 5 Reducing conservatism of DR-ARO via intersection of Wasserstein balls
The proof is in Appendix B.8. For $\lambda=\frac{N}{N+\widehat{N}}$, if the intersection $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$ is nonempty, Proposition 5.8 implies that a sufficient condition for this intersection to include the mixture $\mathbb{Q}_{\mathrm{mix}}$ is $\widehat{\varepsilon}/\varepsilon=N/\widehat{N}$, which is intuitive since the radii of the Wasserstein ambiguity sets are typically chosen inversely proportional to the number of samples [34, Theorem 18].
6 Selecting Wasserstein radii
Our analyses thus far have assumed knowledge of DRO ball radii $\varepsilon$ and $\widehat{\varepsilon}$ (Assumptions 4.2 and 5.1). These are unrealistic in most real-world scenarios. Here we discuss how to set $\varepsilon$ and $\widehat{\varepsilon}$ based on the data such that Problems DR-ARO and Inter-ARO remain well-defined. We consider two settings. First we discuss the case where $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ is known. Then, we discuss the most realistic scenario where this distance is unknown. To this end, we investigate the statistical properties of our distributionally and adversarially robust optimization models to be able to set $\varepsilon$ and $\widehat{\varepsilon}$ values.
Choosing $\epsilon$ in DR-ARO.
@@ 6 Selecting Wasserstein radii
In order to relax Assumption 4.2 in Problem DR-ARO, one needs to infer $\varepsilon$ value from the empirical data so that $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ with a pre-specified level of confidence. The following theorem presents tight characterizations for $\varepsilon$ so that the ball $\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ includes the true distribution $\mathbb{P}^{0}$ with arbitrarily high confidence, and shows that for an $\varepsilon$ chosen in such manner, Problem DR-ARO is well-defined. The detailed statement and the proof are in Appendix B.9.
Theorem 6.1 (abridged).
@@
For light-tailed distribution $\mathbb{P}^{0}$ and $\varepsilon\geq\mathcal{O}(\frac{\log(\eta^{-1})}{N})^{1/n}$ for $\eta\in(0,1)$, we have: (i) $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ with $1-\eta$ confidence; (ii) DR-ARO overestimates true loss with $1-\eta$ confidence; (iii) DR-ARO is asymptotically consistent $\mathbb{P}^{0}$-a.s.; (iv) worst-case distributions for optimal solutions of DR-ARO are supported on at most $N+1$ outcomes.
Choosing $\epsilon$ and $\epsilon^{\prime}$ in Inter-ARO.
Inter-ARO revises DR-ARO by intersecting $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ with another ball $\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$ centered at the auxiliary distribution. We need a nonempty intersection for Inter-ARO to be well-defined. A sufficient condition follows from the triangle inequality: $\varepsilon+\widehat{\varepsilon}\geq\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$. Moreover, provided that $\varepsilon\geq\mathrm{W}(\mathbb{P}_{N},\mathbb{P}^{0})$, a sufficient condition for $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})=\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})$ is $\widehat{\varepsilon}\geq\varepsilon+\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}})+\mathrm{W}(\widehat{\mathbb{P}}_{\widehat{N}},\widehat{\mathbb{P}})$ (cf. Figure 1). While choosing such $\widehat{\varepsilon}$ to reduce the size of the ambiguity set of DR-ARO, we want this intersection to include $\mathbb{P}^{0}$, assuming $\varepsilon$ is set in light of Theorem 6.1. The auxiliary data ^$\widehat{\mathbb{P}}_{\widehat{N}}$ is constructed from instances that are independently sampled from $\widehat{\mathbb{P}}$ and thus Wasserstein finite sample statistics can estimate $\mathrm{W}(\widehat{\mathbb{P}},\widehat{\mathbb{P}}_{\widehat{N}})$. To have confidence guarantees on $\mathbb{P}^{0}\in\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$, however, we must additionally know $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ which we use in the following result. Full statement of the theorem and its proof are in Appendix B.10.
Theorem 6.2 (abridged).
@@
Remark 6.3.
Inter-ARO is not asymptotically consistent, given that $\widehat{N}\rightarrow\infty$ will let $\widehat{\varepsilon}\rightarrow\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ due to the non-zero distance between the true distribution $\mathbb{P}^{0}$ and the auxiliary distribution $\widehat{\mathbb{P}}$. Inter-ARO is thus not useful in asymptotic data regimes, which is not surprising given that we introduced it to reduce the conservatism of DR-ARO which by design arises in non-asymptotic settings.
@@
In the above results, we assumed that $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ is known. However, this is challenging in most practical settings [48] and we estimate it via cross validation (as in the transfer learning and domain adaptation literature, [83]). For some special cases, we can use domain knowledge (e.g., the “Uber vs Lyft” example of [68]). For example, in a differential privacy context, a data holder shares a subset of opt-in data to form $\mathbb{P}_{N}$, and generates a privacy-preserving synthetic dataset from the rest. Due to challenges in synthetic data generation under privacy constraints, the synthetic distribution approximates the true distribution, resulting in a nonzero Wasserstein distance [24, 71]. Using this distance will complete the above discussion. Another research direction relies on $\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}})$ when it is known, especially when synthetic data generators are trained on the empirical dataset. By employing Wasserstein GANs, which minimize the Wasserstein-1 distance, the distance between the generated distribution and the training distribution is minimized. This ensures that the synthetic distribution remains within a small radius of the training distribution [1].
@@ 7 Experiments
We conduct a series of experiments to test the proposed DR ARO models using empirical and auxiliary datasets. We use the following abbreviations: ERM and ARO stand for solving problems 4 (i.e., minimization of the empirical logistic loss) and 8 (i.e., adversarial training for logistic loss), respectively. ARO+Aux refers to solving problem (14), that is, replacing the empirical distribution of 8 with its mixture with auxiliary data. DRO+ARO is solving DR-ARO, which is the Wasserstein DR counterpart of 8. Finally, DRO+ARO+Aux refers to solving Inter-ARO, which revises DR-ARO by intersecting its ambiguity set with a Wasserstein ball built using auxiliary data. Note that, ERM, ARO, and DRO+ARO are oblivious to auxiliary data. Finally, recall that DRO+ARO and DRO+ARO+Aux are the models that we propose. All Wasserstein radii of DR models, as well as the weight parameters of ARO+Aux are cross-validated. Implementation details are in Appendix D.
7.1 UCI datasets (auxiliary data is synthetic)
We compare the out-of-sample error rates of each method on 5 UCI datasets for classification [22, 39].
For each dataset, we run 10 simulations as follows: (i) Select $40\%$ of the data as a test set ($N_{\mathrm{te}}\propto 0.4$); (ii) Sample $25\%$ of the remaining to form a training set ($N\propto 0.6\cdot 0.25$); (iii) The rest ($\widehat{N}\propto 0.6\cdot 0.75$) is used to fit a synthetic generator Gaussian Copula from the SDV package [46], to sample auxiliary data from. The mean errors on the test set are reported in Table 2 for $\ell_{2}$-attacks of strength $\alpha=0.05$. The best error is always achieved by DRO+ARO+Aux, followed by DRO+ARO, DRO+Aux, ARO, ERM, respectively. In Appendix D.1, we report similar results for 5 more UCI datasets along with attack strengths $\alpha\in\{0,0.05,0.2\}$, and share data preprocessing details and standard deviations.
@@ 7.2 MNIST/EMNIST datasets (auxiliary data is out-of-domain)
We use the MNIST [35] digits dataset to classify whether a digit is 1 or 7. For an auxiliary dataset, we use the EMNIST [17] digits dataset, as the authors of [17] summarize that the EMNIST dataset has additional samples “collected from high school students and pose a more challenging problems”. Since EMNIST digits include MNIST digits, we removed the latter from the EMNIST dataset. We simulated the following $20$ times: (i) Sample $1,000$ instances from the MNIST dataset as a training set; (ii) The remaining instances in the MNIST dataset are our test set; (iii) Sample $1,000$ instances from the EMNIST dataset as an auxiliary dataset. Table 3 reports the mean test errors under various adversarial attack regimes. The results are analogous to UCI experiments.
@@ 7.3 Artificial experiments (auxiliary data is perturbed)
We generate empirical and auxiliary datasets by controlling their data-generating distributions in line with the standard practice (more details in Appendix D.3). We simulate $25$ cases, each with $N=100$ training, $\widehat{N}=200$ auxiliary, and $N_{\mathrm{te}}=10,000$ test instances and $n=100$. The performance of benchmark models with varying $\ell_{2}$-attacks is available in Figure 2 (left). ERM provides the worst performance, followed by ARO, and our DRO+ARO+Aux model gives the best performance. The relationship between DRO+ARO and ARO+Aux is not monotonic: the latter works better in larger attack regimes, conforming to the robust overfitting phenomenon. Finally, Adv+DRO+Aux always performs the best. We conduct a similar simulation for datasets with $n=100$, and gradually increase $N=\widehat{N}$ to report median ($50\%\pm 15\%$ quantiles shaded) runtimes of each method (cf. Figure 2, left). The fastest methods is ARO, followed by ERM, ARO+Aux, DRO+ARO, and DRO+ARO+Aux. The slowest is DRO+ARO+Aux as expected, but the runtime still scales graciously.
@@
8 Conclusions and future work
@@
We formulate the distributionally robust counterpart of adversarially robust logistic regression. Additionally, we demonstrate how to effectively utilize appropriately curated auxiliary data (if available) to mitigate the inherent conservatism of distributional robustness. We illustrate the superiority of the proposed approach in terms of out-of-sample performance and confirm its scalability in practical settings.
@@
It would be natural to extend our results to more loss functions as is typical for theoretical DRO studies stemming from logistic regression. Moreover, the recent breakthroughs in the area of foundation models naturally pose the question of whether the ideas presented in this work apply to these models. For example, [78] uses a pre-trained language model (PLM) to generate synthetic pairs of text sequences and labels which are then used to train a downstream model. It would be interesting to adapt our ideas to the text domain to explore robustness in the presence of two PLMs.
Disclaimer:
This paper was prepared for informational purposes by the CDAO group of JPMorganChase and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. J.P. Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix A Notation
@@
Throughout the paper, bold lower case letters denote vectors, while standard lower case letters are reserved for scalars. A generic data instance is modeled as $(\bm{x},y)\in\Xi:=\mathbb{R}^{n}\times\{-1,+1\}$. For any $p>0$, $\lVert\bm{x}\rVert_{p}$ denotes the rational norm $\left(\sum_{i=1}^{n}\lvert x_{i}\rvert^{p}\right)^{1/p}$ and $\lVert\bm{x}\rVert_{p^{\star}}$ is its dual norm where $\frac{1}{p}+\frac{1}{p^{\star}}=1$ with the convention of $/1+1/\infty=1$ 1. The set of probability distributions supported on $\Xi$ is denoted by $\mathcal{P}(\Xi)$. The Dirac measure supported on $\bm{\xi}$ is denoted by $\delta_{\bm{\xi}}$. The logloss is defined as $\ell_{\bm{\beta}}(\bm{x},y)=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}))$ and its associated univariate loss is $L(z)=\log(1+\exp(-z))$ so that $L(y\cdot\bm{\beta}^{\top}\bm{x})=\ell_{\bm{\beta}}(\bm{x},y)$. The exponential cone is denoted by $\mathcal{K}_{\exp}=\mathrm{cl}(\{\bm{\omega}\in\mathbb{R}^{3}:\omega_{1}\geq\omega_{2}\cdot\exp(\omega_{3}/\omega_{2}),\ \omega_{1}>0,\ \omega_{2}>0\})$ where $\mathrm{cl}$ is the closure operator. The Lipschitz modulus of a univariate function $f$ is defined as $\mathrm{Lip}(f):=\sup_{z,z^{\prime}\in\mathbb{R}}\left\{\frac{|f(z)-f(z^{\prime})|}{|z-z^{\prime}|}:z\neq z\right\}$ whereas its effective domain is $\mathrm{dom}(f)=\{z:f(z)<+\infty\}$. For a function $f:\mathbb{R}^{n}\mapsto\mathbb{R}$, its convex conjugate is $f^{*}(\bm{z})=\sup_{\bm{x}\in\mathbb{R}^{n}}\bm{z}^{\top}\bm{x}-f(\bm{x})$. We reserve $\alpha\geq 0$ for the radii of the norms of adversarial attacks on the features and $\varepsilon\geq 0$ for the radii of distributional ambiguity sets.
Appendix B Proofs
B.1 Proof of Proposition 4.1

### v2

@@ 5 MAIN RESULT
For $\lambda=\frac{N}{N+\widehat{N}}$, if the intersection $\mathfrak{B}_{\varepsilon}({\mathbb{P}}_{N})\cap\mathfrak{B}_{\widehat{\varepsilon}}({\widehat{\mathbb{P}}}_{\widehat{N}})$ is nonempty, Proposition 3 implies that a sufficient condition for this intersection to include $\mathbb{Q}_{\mathrm{mix}}$ is $\widehat{\varepsilon}/\varepsilon=N/\widehat{N}$, which is intuitive since the radii of Wasserstein ambiguity sets are chosen inversely proportional to the number of samples (Kuhn et al.,, 2019, Theorem 18).
6 SETTING WASSERSTEIN RADII
Thus far, we have assumed knowledge of DRO ball radii $\varepsilon$ and $\widehat{\varepsilon}$ that satisfy Assumptions 1 and 2. In this section, we employ Wasserstein finite-sample statistics techniques to estimate these values.
@@ 6 SETTING WASSERSTEIN RADII
Setting $\epsilon$ for DR-ARO. In the following theorem, we present tight characterizations for $\varepsilon$ so that the ball $\mathfrak{B}_{\varepsilon}({{\mathbb{P}}}_{{N}})$ includes the true distribution $\mathbb{P}^{0}$ with arbitrarily high confidence. We show that for an $\varepsilon$ chosen in such a manner, DR-ARO is well-defined. The full description of this result is available in our appendices.
Theorem 2 (abridged).
@@
For light-tailed distribution $\mathbb{P}^{0}$ and $\varepsilon\geq\mathcal{O}(\frac{\log(\eta^{-1})}{N})^{1/n}$ for $\eta\in(0,1)$, we have: (i) $\mathbb{P}^{0}\in\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ with $1-\eta$ confidence; (ii) DR-ARO overestimates the expected loss for $\mathbb{P}^{0}$ with $1-\eta$ confidence; (iii) DR-ARO is asymptotically consistent $\mathbb{P}^{0}$-a.s.; (iv) worst-case distributions for optimal solutions of DR-ARO are supported on at most $N+1$ outcomes.
We next derive an analogous result for Inter-ARO.
Choosing $\epsilon$ and $\epsilon^{\prime}$ in Inter-ARO. Recall that Inter-ARO revises DR-ARO by intersecting $\mathfrak{B}_{\varepsilon}(\mathbb{P}_{N})$ with $\mathfrak{B}_{\widehat{\varepsilon}}(\widehat{\mathbb{P}}_{\widehat{N}})$. We need a nonempty intersection for Inter-ARO to be well-defined. A sufficient condition follows from the triangle inequality: $\varepsilon+\widehat{\varepsilon}\geq\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}}_{\widehat{N}})$. We also want this intersection to include $\mathbb{P}^{0}$ with high confidence, in order to satisfy Assumption 2. We next provide a tight characterization for such $\varepsilon,\widehat{\varepsilon}$. The full description of this result is available in our appendices.
Theorem 3 (abridged).
@@
Remark 1.
Inter-ARO is not asymptotically consistent, given that $\widehat{N}\rightarrow\infty$ will let $\widehat{\varepsilon}\rightarrow\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ due to the non-zero constant distance between the true distribution $\mathbb{P}^{0}$ and the auxiliary distribution $\widehat{\mathbb{P}}$. Inter-ARO is thus not useful in asymptotic data regimes.
@@
In Theorem 3, we use $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ explicitly. This distance, however, is typically unknown, and a common approach is to cross-validate it [^1] Note that, in practice, the distance between the unknown true and auxiliary data-generating distributions is also cross-validated in the transfer learning and domain adaptation literature (Zhong et al.,, 2010).. This would be applicable in our setting thanks to Corollary 2, because the relaxation problem 16 is not forced to learn from the auxiliary data unless it is useful, that is, one can find evidence for the usefulness of the auxiliary data via cross-validation. Moreover, there are several domains where $\mathrm{W}(\mathbb{P}^{0},\widehat{\mathbb{P}})$ is known exactly. For some special cases, we can use direct domain knowledge (e.g., the “Uber vs Lyft” example of Taskesen et al., 2021). A very recent example comes from learning from multi-source data, where $\mathbb{P}^{0}$ is named the target distribution and $\widehat{\mathbb{P}}$ is the source distribution (Rychener et al.,, 2024, §1). Another domain is private data release, where a data holder shares a subset of opt-in data to form $\mathbb{P}_{N}$, and generates a privacy-preserving synthetic dataset from the rest. The (privately generated) synthetic distribution has a known nonzero Wasserstein distance from the true data-generating distribution (Dwork et al.,, 2014; Ullman and Vadhan,, 2020). Alternatively, one can directly rely on $\mathrm{W}(\mathbb{P}_{N},\widehat{\mathbb{P}})$ when it is known, especially when synthetic data generators are trained on the empirical dataset. By employing Wasserstein GANs, which minimize the Wasserstein-1 distance, the distance between the generated distribution and the training distribution is minimized. This ensures that the synthetic distribution remains within a small radius of the training distribution (Arjovsky et al.,, 2017).
@@ 7 EXPERIMENTS
We conduct a series of experiments, each having a different source of auxiliary data, to test the proposed DR ARO models using empirical and auxiliary datasets. We use the following abbreviations, where ‘solution’ refers to the optimal $\bm{\beta}$ to make decisions:
    ERM: Solution of problem 4 (i.e., naiv̈e LR);
    ARO: Solution of problem 8 (i.e., adversarially robust LR);
    ARO+Aux: Solution of problem (19) (i.e., replacing the empirical distribution of 8 with its mixture with auxiliary data);
    DRO+ARO: Solution of DR-ARO (i.e., the Wasserstein DR counterpart of 8);
    DRO+ARO+Aux: Solution of 16 (i.e., relaxation of Inter-ARO that revises DR-ARO and intersects its ambiguity set with a Wasserstein ball built using auxiliary data);
Recall that DRO+ARO and DRO+ARO+Aux are the DR models that we propose. Note also that, ERM, ARO, and DRO+ARO are oblivious to auxiliary data. All Wasserstein radii of DR models, and the weight parameters of ARO+Aux are cross-validated from the same grids. All experiments are conducted in Julia (Bezanson et al.,, 2014) and executed on Intel Xeon 2.66GHz processors with 16GB memory in single-core mode. We use MOSEK’s exponential cone optimizer to solve all problems. Implementation details are in the appendices.
7.1 UCI Datasets (Auxiliary Data is Synthetically Generated)
We compare the out-of-sample error rates of each method on 10 UCI datasets for binary classification (Kelly et al.,, 2023). For each dataset, we run 10 simulations as follows: (i) Select $40\%$ of the data as a test set ($N_{\mathrm{te}}\propto 0.4$); (ii) Sample $25\%$ of the remaining to form a training set ($N\propto 0.6\cdot 0.25$); (iii) The rest ($\widehat{N}\propto 0.6\cdot 0.75$) is used to fit a synthetic generator Gaussian Copula from the SDV package (Patki et al.,, 2016), which is then used to generate auxiliary data. The mean errors on the test set are reported in Table 1 for $\ell_{2}$-attacks of strength $\alpha=0.05$. The best error is always achieved by DRO+ARO+Aux, followed by DRO+ARO, DRO+Aux, ARO, ERM, respectively. In our appendices, we report similar results for attack strengths $\alpha\in\{0,0.05,0.2\}$, and share data preprocessing details and standard deviations of out-of-sample errors.
@@ 7.2 MNIST/EMNIST Datasets (Auxiliary Data is Out-of-Domain)
We use the MNIST digits dataset (LeCun et al.,, 1998) to classify whether a digit is 1 or 7. For an auxiliary dataset, we use the larger EMNIST digits dataset (Cohen et al.,, 2017), whose authors summarize that this dataset has additional samples collected from a different group of individuals (high school students). Since EMNIST digits include MNIST digits, we remove the latter from the EMNIST dataset. We simulate the following 25 times: (i) Sample 1,000 instances from the MNIST dataset as a training set; (ii) The remaining instances in the MNIST dataset are our test set; (iii) Sample 1,000 instances from the EMNIST dataset as an auxiliary dataset. Table 2 reports the mean out-of-sample errors in various adversarial attack regimes. The results are analogous to the UCI experiments.
@@ 7.3 Artificial Experiments (Auxiliary Data is Perturbed)
We generate empirical and auxiliary datasets by controlling their data-generating distributions (more details in the appendices). We simulate $25$ cases, each with $N=100$ training, $\widehat{N}=200$ auxiliary, and $N_{\mathrm{te}}=10,000$ test instances and $n=100$ features. The performance of benchmark models with varying $\ell_{2}$-attacks is available in Figure 2 (left). ERM provides the worst performance, followed by ARO. The relationship between DRO+ARO and ARO+Aux is not monotonic: the latter works better in larger attack regimes, conforming to the robust overfitting phenomenon. Finally, Adv+DRO+Aux always performs the best. We conduct a similar simulation for datasets with $n=100$, and gradually increase $N=\widehat{N}$ to report median ($50\%\pm 15\%$ quantiles shaded) runtimes of each method (cf. Figure 2, right). The fastest methods is ARO, followed by ERM, ARO+Aux, DRO+ARO, and DRO+ARO+Aux. The slowest is DRO+ARO+Aux, but the runtime scales graciously.
8 CONCLUSIONS
@@
We formulate the distributionally robust counterpart of adversarially robust logistic regression. Additionally, we demonstrate how to effectively utilize appropriately curated auxiliary data by intersecting Wasserstein balls. We illustrate the superiority of the proposed approach in terms of out-of-sample performance and confirm its scalability in practical settings.
@@
It would be natural to extend our results to more loss functions as is typical for theoretical DRO studies stemming from logistic regression. Moreover, recent breakthroughs in the area of foundation models naturally pose the question of whether our ideas in this work apply to these models. For example, Ye et al., (2022) use pre-trained language model (PLM) to generate synthetic pairs of text sequences and labels which are then used to train downstream models. It would be interesting to adapt our ideas to the text domain to explore robustness in the presence of two PLMs.
Supplementary Materials
9 NOTATION
@@
Throughout the paper, bold lower case letters denote vectors, while standard lower case letters are reserved for scalars. A generic data instance is modeled as $\bm{\xi}=(\bm{x},y)\in\Xi:=\mathbb{R}^{n}\times\{-1,+1\}$. For any $p>0$, $\lVert\bm{x}\rVert_{p}$ denotes the rational norm $\left(\sum_{i=1}^{n}\lvert x_{i}\rvert^{p}\right)^{1/p}$ and $\lVert\bm{x}\rVert_{p^{\star}}$ is its dual norm where $11/p+1/p^{\star}=$ 1 with the convention of $/1+1/\infty=1$ 1. The set of probability distributions supported on $\Xi$ is denoted by $\mathcal{P}(\Xi)$. The Dirac measure supported on $\bm{\xi}$ is denoted by $\delta_{\bm{\xi}}$. The logloss is defined as $\ell_{\bm{\beta}}(\bm{x},y)=\log(1+\exp(-y\cdot\bm{\beta}^{\top}\bm{x}))$ and its associated univariate loss is $L(z)=\log(1+\exp(-z))$ so that $L(y\cdot\bm{\beta}^{\top}\bm{x})=\ell_{\bm{\beta}}(\bm{x},y)$. The exponential cone is denoted by $\mathcal{K}_{\exp}=\mathrm{cl}(\{\bm{\omega}\in\mathbb{R}^{3}:\omega_{1}\geq\omega_{2}\cdot\exp(\omega_{3}/\omega_{2}),\ \omega_{1}>0,\ \omega_{2}>0\})$ where $\mathrm{cl}$ is the closure operator. The Lipschitz modulus of a univariate function $f$ is defined as $\mathrm{Lip}(f):=\sup_{z,z^{\prime}\in\mathbb{R}}\left\{\lvert f(z)-f(z^{\prime})\rvert/\lvert z-z^{\prime}\rvert:z\neq z\right\}$ whereas its effective domain is $\mathrm{dom}(f)=\{z:f(z)<+\infty\}$. For a function $f:\mathbb{R}^{n}\mapsto\mathbb{R}$, its convex conjugate is $f^{*}(\bm{z})=\sup_{\bm{x}\in\mathbb{R}^{n}}\bm{z}^{\top}\bm{x}-f(\bm{x})$. We reserve $\alpha\geq 0$ for the radii of the norms of adversarial attacks on the features and $\varepsilon\geq 0$ for the radii of distributional ambiguity sets.
10 MISSING PROOFS
10.1 Proof of Observation 1